The velocity of electron in ground state of 
hydrogen atom is
 

Grade:12th pass

3 Answers

Naveen Kumar
9 years ago
angular momentum
mvr=nh/2π
v=nh/2mπr={nh/2πm}* 1/(n2h2/2πmKe2)=2.2*106m/s
where
n = 1
K = 9 x 109Nm2/c2
m = 9.1 x 10-31kg
e = 1.6 x 10-19c,
r = 5.29 x 10-11m
Shobhit Varshney IIT Roorkee
9 years ago
Hi,

use Coulombic force between e and nucleus= centripetal force on electron

thanks.....................................................................................................................................................................................

sajid
8 years ago
Nm2/c2m = 9.1 x 10-31kge = 1.6 x 10-19c,r = 5.29 x 10-11m9m/swheren = 1K = 9 x 106=2.2*102)/2πmKe2h2angular momentummvr=nh/2πv=nh/2mπr={nh/2πm}* 1/(n